solve by substitution indicate whether each systm is independent inconsistent or dependent
-3x+y=3
2x-3y=5

Answer by mukhopadhyay(490)

-3x+y=3
=>y=3+3x
2x-3y=5
Substitute the value of y in the above equation.
2x-3(3+3x)=5
=>2x-9-9x=5
=>-7x=5-9
=>-7x=14
=>x=-2
Since y=3+3x, y=3+3(-2)=3-6=-3
Answer: x=-2 and y=-3